OpenStudy (anonymous):

Given F(x) shown below, complete the equation for the inverse of F(x). If necessary, use the slash mark (/) for the division symbol.

OpenStudy (anonymous):

\[F(x)=9x-8\]

OpenStudy (anonymous):

@Anickyan @Luigi0210 help me please

OpenStudy (anonymous):

would it be \[x+9/8\]

OpenStudy (luigi0210):

Okay, now I'm certain that for this one you switch y and x

OpenStudy (luigi0210):

Close, you just forgot to apply the division to the x

OpenStudy (anonymous):

how do i do that

OpenStudy (anonymous):

Are you familiar with solving equations?

OpenStudy (anonymous):

I can tell you what you did wrong: instead of dividing both x and 8 by 9, you divided 9 by 8. The inverse function should look like this: \[f⁻¹(x) = \frac{x - 8}{9}\]

OpenStudy (anonymous):

oh so thats how my answer should look like?

OpenStudy (anonymous):

Yes. You could also replace \(x\) with \(y\), it wouldn't matter.

OpenStudy (anonymous):

not exactly... y = 9x - 8 now switch x and y to get x = 9y - 8 and solve for y... you should get y = (x+8)/9

OpenStudy (anonymous):

Yeah sorry, what @pgpilot326 , I solved with \(9x + 8\) instead of \(9x - 8\) My bad! :)

OpenStudy (anonymous):

no worries!!! you had the right idea!!!

OpenStudy (luigi0210):

Yup yup..